# EvoP: Robust LLM Inference via Evolutionary Pruning

### v1

@@ 6 Related Work
Language Models Compression. There are several main techniques for language model compression: knowledge distillation Sun et al. (2020); Touvron et al. (2021); Pan et al. (2021), quantization Yao et al. (2022); Gholami et al. (2021); Xiao et al. (2023), network pruning or sparsity, He et al. (2024); Frantar and Alistarh (2023); Ashkboos et al. (2024); Sun et al. (2024); Zhang et al. (2024b); Song et al. (2024); Liu et al. (2023) and early exit Huang et al. (2024); Li et al. (2023); Xin et al. (2020). Knowledge distillation methods transfer knowledge from a large, complex model (called the teacher model) to a smaller, simpler model (called the student model). They must either learn the output distribution of teacher models Agarwal et al. (2024) or design multi-task approaches Liang et al. (2023) to ensure that student models retain the knowledge and generative capabilities of the teacher models. Quantization methods compress language models by reducing the precision of the numerical values representing the model’s parameters (e.g., weights and activations). For example, OneBit quantized the weight matrices of LLMs to 1-bit Xu et al. (2024). Early exit methods allow a model to terminate its processing early during inference if it has already made a confident prediction, avoiding the need for additional layers of computation Chen et al. (2024). Network pruning, also known as sparsity techniques, refers to methods employed to compress language models by eliminating less significant structures within the network, such as individual weights, neurons, or layers Frantar and Alistarh (2023); Ashkboos et al. (2024); Song et al. (2024). The primary objectives are to reduce the model’s size, enhance inference speed, and decrease memory consumption while preserving or only marginally affecting its performance. These works are orthogonal to each other, and we focus on the pruning methods.
Different Granularity of Pruning. Recent studies have investigated pruning at various granularities, ranging from coarse to fine. At the coarse-grained level, pruning methods include layer-wise Song et al. (2024), attention heads Voita et al. (2019), channel-wise pruning Ashkboos et al. (2024); Ma et al. (2023), and neurons Jiang et al. (2018). At the fine-grained level, techniques such as N:M sparsity Frantar and Alistarh (2023); Sun et al. (2024); Zhang et al. (2024b); Zhou et al. (2021); Zhang et al. (2022b) and individual weight pruning Wang (2020) have been explored. This paper mainly explored but was not limited to layer-wise network pruning.
@@ 7 Conclusion
8 Limitations
The search time is the main limitation of the proposed EvoP, which is longer than existing pruning methods and usually takes about hours to run on one NVIDIA A100 GPU. However, this search time should be accepted. The reasons are: 1) This is an offline process and only requires running once for multiple deployments; 2) With the searched pruning pattern, the pruned model can consistently improve its performance, making it worthwhile to spend the time cost; 3) Furthermore, parallel or distributed computing can accelerate the evolutionary pruning pattern searching process, while existing implementations lack parallelism.

### v2

@@ 6 Related Work
Network pruning, also known as sparsity techniques, refers to methods employed to compress language models by eliminating less significant structures within the network, such as individual weights, neurons, or layers [4, 1, 13]. The primary objectives are to reduce the model’s size, enhance inference speed, and decrease memory consumption while preserving or only marginally affecting its performance. Recent studies have investigated pruning at various granularities, ranging from coarse to fine. At the coarse-grained level, pruning methods include layer-wise [13], attention heads [18], channel-wise pruning [1, 9], and neurons [7]. At the fine-grained level, techniques such as N:M sparsity [4, 14, 25, 26, 24] and individual weight pruning [19] have been explored. This paper mainly explored but was not limited to layer-wise network pruning.